# Unveiling Super Experts in Mixture-of-Experts Large Language Models

### v2

@@ 1 Introduction
Sparsely activated Mixture-of-Experts (MoE) models employ dynamic routing and sparse activation, demonstrating significant potential in enhancing the learning capacity of large language models (LLMs) (Cai et al., 2024; Mu & Lin, 2025). This paradigm has led to the development of state-of-the-art MoE LLMs, including DeepSeek (Guo et al., 2025; Liu et al., 2024b), Qwen (Yang et al., 2025a), LongCat-Flash (Team et al., 2025) and others. Despite their potential, a significant challenge stems from their large parameter size and high computational cost (Li et al., 2023; Lu et al., 2024; Chowdhury et al., 2024), which present considerable obstacles for deployment. Model compression techniques, such as quantization (Frantar et al., 2022; Xiao et al., 2023a; Su et al., 2025a), pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and others (Zhu et al., 2024; Wang et al., 2024), enable the development of more compact and computationally efficient models.
Beyond LLM-oriented compression approaches, expert-level compression methods have been developed by leveraging the structural characteristics of MoE models and the uneven importance of experts induced by training strategies (Chowdhury et al., 2024; Chi et al., 2022; Lu et al., 2024). Specifically, it employs various expert importance metrics to guide the pruning, merging, or skipping of less critical experts (Lu et al., 2024; Huang et al., 2025; Xie et al., 2024), prioritize more important ones by assigning higher bit budgets during quantization (Duanmu et al., 2025; Li et al., 2024), and allocate more ranks in low-rank decomposition (Yang et al., 2024a; Li et al., 2023).
@@
For instance, several works evaluate expert importance by measuring activation frequency or by analyzing router scores produced within MoE layers (Li et al., 2024; 2023; Huang et al., 2025). Additionally, reconstruction loss and other similarity-based metrics have been utilized in previous studies (Lu et al., 2024; Duanmu et al., 2025; Zhang et al., 2024b).
Analyzing expert importance not only facilitates model compression but also provides deeper insights into the inner workings of MoE LLMs. However, existing approaches often rely on empirical criteria to identify critical experts, lacking a deeper exploration and understanding of the heterogeneous importance among experts. In this study, we address a fundamental yet previously overlooked question: Is there a small subset of distinct experts that plays an exceptionally critical role in the underlying mechanisms of MoE LLMs?
@@ 3.2.1 Super Experts Profiling
where $P_{99.5}=\text{Percentile}_{99.5}(\mathcal{A})$ and $a_{\text{max}}=\max\mathcal{A}$. This criterion effectively identifies the experts of interest across various MoE LLMs we investigated, as highlighted in bold in Table 1. No specific dataset is designated for identifying SEs, since we later demonstrate that their distribution remains stable across different input datasets. The pseudocode of SEs profiling is presented in Appendix I. We have developed an automated tool for rapid and precise SE profiling based on this definition. The code is provided in https://github.com/ZunhaiSu/Super-Experts-Profilling.
@@ 3.2.2 Distribution of Super Experts Across Models and Data Domains
Qwen3-30B-A3B
Layer 1 Expert 68, Layer 2 Expert 92,
Layer 3 Expert 82
@@
For instance, Qwen3 and Mixtral do not employ shared experts, whereas DeepSeek does. DeepSeek-V2-Lite adopts a hybrid architecture, wherein the first layer utilizes dense MLPs, while the remaining layers are based on MoE blocks. Through the proposed SE profiling tool, we identify the SEs in these models using the C4 (Raffel et al., 2020) dataset. A summary of the SEs is provided in Table 2, and heatmap visualizations of the maximum output magnitudes from the down_proj are shown in Figure 5. The key conclusions regarding SEs are summarized as follows: (i) SEs are consistently present across the investigated models, accounting for less than 0.5% of all experts. (ii) After post-training processes, the distribution of SEs remains unchanged compared to the base model. Moreover, some experts in the final layers also exhibit extreme activation outliers. However, since they do not contribute to the formation of MAs, they do not hold the same level of significance as SEs. Additional results are available in Appendix C.
@@
DeepSeek
V2-Lite
Baseline 60.27 46.59 78.37 79.79 37.83 58.75 55.03 34.60 80.30 71.19 6.31
@@
Model Setting Avg. GPQA Diamond Math-500 AIME 2024 AIME 2025 LiveCodeBench
Wiki
PPL
@@
Model Setting Avg. GPQA Diamond Math-500 AIME 2024 AIME 2025 HumanEval
Wiki
PPL
@@ 4.2 Impact on Reasoning Models
The results, presented in Tables 4 and 5, show that pruning the SEs causes a significant performance degradation, while random pruning has almost no impact. The Pass@1 scores for most tasks drop to zero, highlighting the critical role of SEs. During the review of model responses on the Math-500 benchmark, we made a striking observation: after pruning the SEs, the model consistently generated repetitive responses in nearly every test, continuing until it reached the maximum output length, as shown in Table 9 and 10. This behavior suggests that the model loses its ability to reason and solve problems entirely after SE pruning, with additional discussion on this part provided in Appendix H.
@@ 5.1 Super Experts as the Origin of Systematic Outliers in MoE LLMs
Building on prior research and our own findings on SEs, we demonstrate that SEs constitute the fundamental source of systematic outliers in MoE LLMs. Specifically, using Qwen3-10B-A3B as an example, the router scores assigned to SEs for the first token (which also serves as the attention sink token) are exceptionally large, whereas for non-sink tokens the scores are more evenly distributed across experts, as shown in the visualization calibrated on the C4 dataset in Figure 6. This routing behavior of SEs ensures that the attention sink token is strongly activated at the SEs. Notably, this behavior is independent of the input dataset (see Appendix F), which also explains why SEs distributions are model-specific. The sink token subsequently produces activation outliers in the output through weight outliers in the down_proj. Through the residual connections, these outliers propagate into the hidden states as MAs. At the attention layers, such tokens then attract disproportionate attention and ultimately emerge as attention sinks. Unlike dense LLMs, where such behavior typically occurs within a single layer (An et al., 2025; Su & Yuan, 2025), MoE models exhibit the progressive formation of systematic outliers by SEs across multiple layers. The overall process is illustrated in Figure 7. More detailed analyses of this process are presented in Appendix G.
@@ Appendix C Further Analysis of Outlier Experts in Final Layers
(ii) In both Qwen3-30B-A3B and DeepSeek-R1, pruning outlier experts does not result in repetitive outputs on reasoning benchmarks such as Math-500 (Lightman et al., 2023), whereas pruning SEs does, as shown in Tables 9 and 10. The pruned SEs and outlier experts is shown in Table 7.
(iii) We observed on Qwen3-30B-A3B that the distribution of outlier experts varies with the input dataset, while the distribution of SEs remains quite stable, as illustrated in Figure 11 and 12.
@@ Appendix D Distribution of Super Experts Across Various Data Domains
In addition to analyzing the distribution of SEs across different models based on the C4 dataset (Raffel et al., 2020), we also examine their distribution patterns across various input data domains. We assess the impact of diverse language inputs on SEs using the WikiText-2 (Merity et al., 2016) and C-Eval (Huang et al., 2023) datasets. Furthermore, we investigate the influence of data from the mathematics and coding domains using the GSM8K (Hendrycks et al., 2021) and HumanEval (Chen et al., 2021) datasets. As shown in Figures 11, 12, 13, 14, 15, and 16, the distribution of SEs remains highly stable, regardless of variations in the input data domain.
@@ Appendix E Additional Results of Reasoning Models After Super Experts Pruning
After pruning SEs, we consistently observed repetitive output and a loss of reasoning ability in both Qwen3-30B-A3B and DeepSeek-R1. The pruned SEs are shown in Table 7, and additional examples from the Math-500 (Lightman et al., 2023) benchmark are presented in Tables 9 and 10.
@@ Appendix H Weight-level Analyses of Super Experts
These SWs are subsequently pruned during inference, and experiments are conducted on LLaVA-V1.5-7B (Liu et al., 2023), Llama3.2-1B (Touvron et al., 2023), and Qwen3-30B-A3B, spanning dense LLMs, vision-language models (VLMs), and MoE LLMs. Tables 11, 12, 13, and 14 show that removing the SWs consistently leads to repetitive and uninformative outputs across all models, demonstrating that these weights are critical contributors to SE importance. This outcome is expected, as it is consistent with our analysis of SEs’ role in driving systematic outliers in Transformers. Interestingly, in dense LLMs, the single FFN layer exhibiting massive activations can be seen as playing a role analogous to that of an SE.
Why does the model exhibit a substantial performance gap before and after SWs pruning? We then investigate the underlying causes of this discrepancy. As model neurons are often polysemantic (Gao et al., 2024a), directly analyzing the semantics of SWs is challenging due to their intrinsic polysemanticity. To address this, we employ Sparse Autoencoders (SAEs) Bloom et al. (2024); Gao et al. (2024a), an unsupervised method designed to extract interpretable features from LLMs by reconstructing activations through a sparse bottleneck layer. Using SAEs, we can decompose the semantics of polysemantic neurons into more discrete, monosemantic features and conduct semantic analyses on the features most strongly correlated with SWs. We then train our SAE using activations from the Layer 1 FFN outputs of Llama-3.2-1B with the C4 dataset (Raffel et al., 2020). The loss function consists of two components: a reconstruction loss and a sparsity penalty loss (Bloom et al., 2024). The reconstruction loss is defined as
@@
Appendix I Algorithm for Profiling Super Experts

### v3

@@ 1 Introduction
Sparsely activated Mixture-of-Experts (MoE) models employ dynamic routing and sparse activation, demonstrating significant potential in enhancing the learning capacity of large language models (LLMs) (Cai et al., 2024; Mu & Lin, 2025). This paradigm has led to the development of state-of-the-art MoE LLMs, including DeepSeek (Guo et al., 2025; Liu et al., 2024b), Qwen (Yang et al., 2025a), LongCat-Flash (Team et al., 2025b; a) and others. Despite their potential, a significant challenge stems from their large parameter size and high computational cost (Li et al., 2023; Lu et al., 2024; Chowdhury et al., 2024), which present considerable obstacles for deployment. Model compression techniques, such as quantization (Frantar et al., 2022; Xiao et al., 2023a; Su et al., 2025a), pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and others (Zhu et al., 2024; Wang et al., 2024), enable the development of more compact and computationally efficient models.
Beyond LLM-oriented compression approaches, expert-level compression methods have been developed by leveraging the structural characteristics of MoE models and the uneven importance of experts induced by training strategies (Chowdhury et al., 2024; Chi et al., 2022; Lu et al., 2024).
@@
Specifically, it employs various expert importance metrics to guide the pruning, merging, or skipping of less critical experts (Lu et al., 2024; Huang et al., 2025; Xie et al., 2024), prioritize more important ones by assigning higher bit budgets during quantization (Duanmu et al., 2025; Li et al., 2024), and allocate more ranks in low-rank decomposition (Yang et al., 2024a; Li et al., 2023). For instance, several works evaluate expert importance by measuring activation frequency or by analyzing router scores produced within MoE layers (Li et al., 2024; 2023; Huang et al., 2025). Additionally, reconstruction loss and other similarity-based metrics have been utilized in previous studies (Lu et al., 2024; Duanmu et al., 2025; Zhang et al., 2024b).
Analyzing expert importance not only facilitates model compression but also provides deeper insights into the inner workings of MoE LLMs Zhang et al. (2026). However, existing approaches often rely on empirical criteria to identify critical experts, lacking a deeper exploration and understanding of the heterogeneous importance among experts. In this study, we address a fundamental yet previously overlooked question: Is there a small subset of distinct experts that plays an exceptionally critical role in the underlying mechanisms of MoE LLMs?
@@ 3.2.1 Super Experts Profiling
where $P_{99.5}=\text{Percentile}_{99.5}(\mathcal{A})$ and $a_{\text{max}}=\max\mathcal{A}$. This criterion is motivated by the heavy-tailed distribution of $a_{l,e}$ and effectively identifies the experts of interest across various MoE LLMs, as highlighted in bold in Table 1. Additional analyses are provided in Appendix I. No specific dataset is designated for identifying SEs, since we later demonstrate that their distribution remains stable across different input datasets. The pseudocode of SEs profiling is presented in Appendix L. We have developed an automated tool for rapid and precise SE profiling based on this definition. The code is provided in https://github.com/ZunhaiSu/Super-Experts-Profilling.
@@ 3.2.2 Distribution of Super Experts Across Models and Data Domains
Qwen3-30B-A3B Layer 1 Expert 68, Layer 2 Expert 92, Layer 3 Expert 82
@@
For instance, Qwen3 and Mixtral do not employ shared experts, whereas DeepSeek does. DeepSeek-V2-Lite adopts a hybrid architecture, wherein the first layer utilizes dense MLPs, while the remaining layers are based on MoE blocks. Through the proposed SE profiling tool, we identify the SEs in these models using the C4 (Raffel et al., 2020) dataset. A summary of the SEs is provided in Table 2, and heatmap visualizations of the maximum output magnitudes from the down_proj are shown in Figure 5. The key conclusions regarding SEs are summarized as follows: (i) SEs are consistently present across the investigated models, accounting for less than 0.5% of all experts. (ii) After post-training processes, the distribution of SEs remains unchanged compared to the base model. Additional results on the distribution of SEs across training stages is provided in Appendix K. Moreover, some experts in the final layers also exhibit extreme activation outliers. However, since they do not contribute to the formation of MAs, they do not hold the same level of significance as SEs. Additional results are available in Appendix C.
@@
DeepSeek V2-Lite Baseline 60.27 46.59 78.37 79.79 37.83 58.75 55.03 34.60 80.30 71.19 6.31
@@
Model Setting Avg. GPQA Diamond Math-500 AIME 2024 AIME 2025 LiveCodeBench Wiki PPL
@@
Model Setting Avg. GPQA Diamond Math-500 AIME 2024 AIME 2025 HumanEval Wiki PPL
@@ 4.2 Impact on Reasoning Models
The results, presented in Tables 4 and 5, show that pruning the SEs causes a significant performance degradation, while random pruning has almost no impact. The Pass@1 scores for most tasks drop to zero, highlighting the critical role of SEs. During the review of model responses on the Math-500 benchmark, we made a striking observation: after pruning the SEs, the model consistently generated repetitive responses in nearly every test, continuing until it reached the maximum output length, as shown in Table 11 and 12. This behavior suggests that the model loses its ability to reason and solve problems entirely after SE pruning, with additional discussion on this part provided in Appendix H. More results are in Appendix E.
@@ 5.1 Super Experts as the Origin of Systematic Outliers in MoE LLMs
Building on prior research and our own findings on SEs, we demonstrate that SEs constitute the fundamental source of systematic outliers in MoE LLMs. Specifically, using Qwen3-10B-A3B as an example, the router scores assigned to SEs for the first token (which also serves as the attention sink token) are exceptionally large, whereas for non-sink tokens the scores are more evenly distributed across experts, as shown in the visualization calibrated on the C4 dataset in Figure 6. This routing behavior of SEs ensures that the attention sink token is strongly activated at the SEs. Notably, this behavior is independent of the input dataset (see Appendix F), which also explains why SEs distributions are model-specific. The sink token subsequently produces activation outliers in the output through weight outliers in the down_proj. Through the residual connections, these outliers propagate into the hidden states as MAs. At the attention layers, such tokens then attract disproportionate attention and ultimately emerge as attention sinks. Unlike dense LLMs, where such behavior typically occurs within a single layer (An et al., 2025; Su & Yuan, 2025), MoE models exhibit the progressive formation of systematic outliers by SEs across multiple layers. The overall process is illustrated in Figure 7. More detailed analyses of this process are presented in Appendix G. Additional cross-domain analyses of SEs and attention sink tokens are provided in Appendix J.
@@ Appendix C Further Analysis of Outlier Experts in Final Layers
(ii) In both Qwen3-30B-A3B and DeepSeek-R1, pruning outlier experts does not result in repetitive outputs on reasoning benchmarks such as Math-500 (Lightman et al., 2023), whereas pruning SEs does, as shown in Tables 11 and 12. The pruned SEs and outlier experts is shown in Table 7.
(iii) We observed on Qwen3-30B-A3B that the distribution of outlier experts varies with the input dataset, while the distribution of SEs remains quite stable, as illustrated in Figure 14 and 15.
(iv) The router scores in the final layer are relatively evenly distributed across all experts for both sink and non-sink tokens, as shown in Figure 10. In contrast, sink-token router logits are strongly skewed toward SEs, indicating that the final-layer outlier experts are fundamentally different from SEs.
(a) Layer 1 sink token.
(b) Layer 1 non-sink tokens.
(c) Layer 2 sink token.
(d) Layer 2 non-sink token.
(e) Layer 3 sink token.
(f) Layer 3 non-sink tokens.
(g) Layer 47 sink token.
(h) Layer 47 non-sink tokens.
@@ Appendix D Distribution of Super Experts Across Various Data Domains
In addition to analyzing the distribution of SEs across different models based on the C4 dataset (Raffel et al., 2020), we also examine their distribution patterns across various input data domains. We assess the impact of diverse language inputs on SEs using the WikiText-2 (Merity et al., 2016) and C-Eval (Huang et al., 2023) datasets. Furthermore, we investigate the influence of data from the mathematics and coding domains using the GSM8K (Hendrycks et al., 2021) and HumanEval (Chen et al., 2021) datasets. As shown in Figures 14, 15, 16, 17, 18, and 19, the distribution of SEs remains highly stable, regardless of variations in the input data domain.
@@ Appendix E Additional Results of Reasoning Models After Super Experts Pruning
After pruning SEs, we consistently observed repetitive output and a loss of reasoning ability in both Qwen3-30B-A3B and DeepSeek-R1. The pruned SEs are shown in Table 7, and additional examples from the Math-500 (Lightman et al., 2023) benchmark are presented in Tables 11 and 12.
@@ Appendix H Weight-level Analyses of Super Experts
These SWs are subsequently pruned during inference, and experiments are conducted on LLaVA-V1.5-7B (Liu et al., 2023), Llama3.2-1B (Touvron et al., 2023), and Qwen3-30B-A3B, spanning dense LLMs, vision-language models (VLMs), and MoE LLMs. Tables 13, 14, 15, and 16 show that removing the SWs consistently leads to repetitive and uninformative outputs across all models, demonstrating that these weights are critical contributors to SE importance. This outcome is expected, as it is consistent with our analysis of SEs’ role in driving systematic outliers in Transformers. Interestingly, in dense LLMs, the single FFN layer exhibiting massive activations can be seen as playing a role analogous to that of an SE.
Why does the model exhibit a substantial performance gap before and after SWs pruning? We then investigate the underlying causes of this discrepancy. As model neurons are often polysemantic (Gao et al., 2024a), directly analyzing the semantics of SWs is challenging due to their intrinsic polysemanticity. To address this, we employ Sparse Autoencoders (SAEs) (Bloom et al., 2024; Gao et al., 2024a), an unsupervised method designed to extract interpretable features from LLMs by reconstructing activations through a sparse bottleneck layer. Using SAEs, we can decompose the semantics of polysemantic neurons into more discrete, monosemantic features and conduct semantic analyses on the features most strongly correlated with SWs. We then train our SAE using activations from the Layer 1 FFN outputs of Llama-3.2-1B with the C4 dataset (Raffel et al., 2020). The loss function consists of two components: a reconstruction loss and a sparsity penalty loss (Bloom et al., 2024). The reconstruction loss is defined as
@@
Appendix I Analysis of Threshold-Based Super Experts Identification
[…]
Table 9: Sensitivity analysis of SE identification across a range of threshold settings for multiple MoE LLMs and input domains.
Models Number of Identified Experts
Qwen3-30B-A3B C4 $P_{95}$ $P_{99}$ $P_{99.5}$ $P_{99.9}$ WikiText-2 $P_{95}$ $P_{99}$ $P_{99.5}$ $P_{99.9}$
$0.1*a_{\text{max}}$ 3 3 3 3 $0.1*a_{\text{max}}$ 3 3 3 3
$0.09*a_{\text{max}}$ 3 3 3 3 $0.09*a_{\text{max}}$ 3 3 3 3
$0.08*a_{\text{max}}$ 3 3 3 3 $0.08*a_{\text{max}}$ 3 3 3 3
$0.07*a_{\text{max}}$ 3 3 3 3 $0.07*a_{\text{max}}$ 3 3 3 3
DeepSeek-V2-Lite C4 $P_{95}$ $P_{99}$ $P_{99.5}$ $P_{99.9}$ WikiText-2 $P_{95}$ $P_{99}$ $P_{99.5}$ $P_{99.9}$
$0.1*a_{\text{max}}$ 2 2 2 2 $0.1*a_{\text{max}}$ 2 2 2 2
$0.09*a_{\text{max}}$ 2 2 2 2 $0.09*a_{\text{max}}$ 2 2 2 2
$0.08*a_{\text{max}}$ 2 2 2 2 $0.08*a_{\text{max}}$ 2 2 2 2
$0.07*a_{\text{max}}$ 2 2 2 2 $0.07*a_{\text{max}}$ 2 2 2 2
Mixtral-8x7B C4 $P_{95}$ $P_{99}$ $P_{99.5}$ $P_{99.9}$ WikiText-2 $P_{95}$ $P_{99}$ $P_{99.5}$ $P_{99.9}$
$0.1*a_{\text{max}}$ 1 1 1 1 $0.1*a_{\text{max}}$ 1 1 1 1
$0.09*a_{\text{max}}$ 1 1 1 1 $0.09*a_{\text{max}}$ 1 1 1 1
$0.08*a_{\text{max}}$ 1 1 1 1 $0.08*a_{\text{max}}$ 1 1 1 1
$0.07*a_{\text{max}}$ 1 1 1 1 $0.07*a_{\text{max}}$ 1 1 1 1
In this section, we demonstrate that SEs reflect an intrinsic property of MoE LLMs, with threshold-based identification serving as a practical and principled detection method. We first provide a justification based on the heavy-tailed distribution of $\mathcal{A}=\{a_{l,e}\}$ (as discussed in Section 3.2.1), and then analyze the robustness of the threshold-based SE identification.
The identification of SEs is motivated by the heavy-tailed nature of extreme activations. While most experts produce modest responses, a small subset consistently generates extreme values (as illustrated in the heatmaps in Figure 5 and Appendix D), forming a distinct long tail that is naturally present in MoE LLMs. To mathematically validate this observation, we analyze the tail of the activation distribution using standard heavy-tail analysis techniques, namely the complementary cumulative distribution function (CCDF) and power-law fitting. The CCDF of a random variable $X$ is defined as
$\text{CCDF}(x)=P(X>x),$ (10)
representing the probability that $X$ exceeds a given value $x$. Plotting the CCDF on a log-log scale provides an initial visual assessment of the tail behavior. To quantitatively characterize the tail, we fit it to a power-law model
$P(X>x)\propto x^{-\alpha},\quad x\geq x_{\min},$ (11)
where $\alpha$ is the tail exponent and $x_{\min}$ denotes the minimum value above which the power-law behavior holds. The exponent $\alpha$ is estimated using maximum likelihood estimation (MLE). As shown in Figure 12, the CCDF tail is consistent with a power-law model. This long-tail behavior indicates that only a few experts dominate the extreme activations, naturally separating them from the majority and confirming that extreme activations follow an inherent heavy-tailed distribution rather than arising from random noise. Threshold-based methods thus provide a practical and principled approach for identifying this small, statistically significant subset of experts.
To further assess the robustness of SE identification, we perform a sensitivity analysis over a reasonable range of thresholds across different MoE LLMs and input domains. As shown in Table 9, SE identification remains consistent, confirming that extreme activation outliers reflect an intrinsic property of MoE LLMs rather than an artifact of any specific threshold choice. Together, these analyses demonstrate that the threshold-based criterion is both robust and scientifically justified, and that SEs constitute an inherent feature of MoE LLMs.
Appendix J Cross-Domain Analysis of Super Experts and Attention Sink Tokens
In this section, we conduct a cross-domain analysis of SEs and attention sink tokens using inputs drawn from diverse domains, including out-of-distribution (OoD) inputs that differ substantially from the training datasets. Each domain includes visualizations of SE heatmaps, attention sink patterns, and router score distributions for both sink and non-sink tokens.
We use C4, C-Eval, GSM8K, and HumanEval (Raffel et al., 2020; Huang et al., 2023; Hendrycks et al., 2021; Chen et al., 2021), covering a range of domains including English, Chinese, mathematics, and code. For the OoD datasets, we adopt Pile-of-Law and MedDialog (Henderson* et al., 2022; Chen et al., 2020), which are legal and medical datasets and differ significantly from the pretraining corpus of the tested MoE LLMs in content and domain.
As shown in Figures 24, 25, 26, 27, 28 and 29, the distribution of SEs remains highly consistent across these varied domains and OoD datasets. Moreover, the sink tokens consistently appear as the first token. Their router scores on SEs are markedly larger, whereas the router scores of non-sink tokens remain relatively uniform across experts. These observations reinforce our claim that SEs reflect an intrinsic property of MoE LLMs, rather than arising from over-exposure to domain-specific tokens, and are not affected by OoD inputs.
Appendix K Distribution of Super Experts Across Training Stages
(a) Qwen3-30B-A3B.
(b) Qwen3-30B-A3B-Base.
(c) Qwen3-30B-A3B-Instruct-2507
(d) Qwen3-30B-A3B-Thinking-2507.
Figure 13: Heatmap visualizations of the maximum output magnitudes from the down_proj for each expert across layers based on C4 dataset. SEs are highlighted with arrows.
Table 10: Super Experts in Qwen3-30B-A3B.
Models Super Experts
Qwen3-30B-A3B-Base Layer 1 Expert 68, Layer2 Expert 92, Layer3 Expert 82
Qwen3-30B-A3B
Qwen3-30B-A3B-Instruct-2507
Qwen3-30B-A3B-Thinking-2507
To study the distribution of SEs across training stages, we identified SEs in four officially open-sourced Qwen3-30B-A3B models at different training stages: Qwen3-30B-A3B-Base, Qwen3-30B-A3B, Qwen3-30B-A3B-Instruct-2507, and Qwen3-30B-A3B-Thinking-2507. The indices of the SEs are shown in Table 10, and the SE heatmaps are presented in Figure 13, demonstrating fully consistent sets across all models and training stages. These results support the conclusion that SEs are persistent features of MoE LLMs.
Appendix L Algorithm for Profiling Super Experts